# Beyond Memorization: Evaluating the True Type Inference Capabilities of LLMs for Java Code Snippets

### v1

@@ Abstract.
Type inference is a crucial task for reusing online code snippets, often found on platforms like StackOverflow, which frequently lack essential type information such as fully qualified names (FQNs) and required libraries. Recent studies have leveraged Large Language Models (LLMs) to perform type inference for such code snippets, demonstrating promising performance. However, these results are potentially affected by a data leakage issue, as the benchmark suite (StatType-SO) used for evaluation has been publicly available on GitHub since 2017, with the full suite available since 2023. Thus it remains uncertain whether the LLMs’ strong performance stems from their ability to understand the semantics of the code snippets or merely from retrieving the ground truth from their training data.
To comprehensively assess the type inference capabilities of LLMs on Java code snippets and identify potential limitations of LLMs, we conducted a three-pronged evaluation. First, utilizing Thalia, a program synthesis technique, we created ThaliaType—a new, previously unreleased dataset designed for type inference evaluation. ThaliaType comprises 300 unseen code snippets, generated in a manner consistent with the setting of StatType-SO. Our results revealed that LLMs exhibit a significant drop in performance when generalizing to unseen code snippets with up to 59% decrease in precision and up to 72% decrease in recall. To further investigate the limitations of LLMs in understanding the semantic meaning of the code, we developed semantic-preserving code transformations. Our analysis showed that LLMs performed significantly worse on code snippets that are syntactically different but semantically equivalent. Finally, we employed delta debugging to distill the minimal syntax elements in snippets sufficient for inference by LLMs. We observe that while the type inference task primarily involves inferring FQNs for types used in code snippets, LLMs could still correctly infer FQNs even when the types were absent from the snippets. This suggests that LLMs rely more on re-using knowledge from training data instead of thoroughly analyzing the snippets.
Our experiments suggest that the strong performance of LLMs in prior evaluations was likely influenced by data leakage in the benchmarks, rather than a genuine understanding of the semantics of code snippets. Our findings highlight the crucial need for carefully designed and rigorously evaluated benchmarks to accurately assess the true capabilities of LLMs for type inference tasks. To ensure reliable and rigorous evaluations of LLMs’ type inference capabilities, future research should prioritize the use of unseen benchmarks that are explicitly excluded from the LLMs’ training data.
@@ 1. Introduction
Type inference for code snippets can assist developers in recovering missing type information needed to reuse these code snippets (Dong et al., 2022; Subramanian et al., 2014; Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023b; Kabir et al., 2024). Existing approaches to type inference for Java code snippets generally fall into two categories: constraint-based and machine learning (ML)-based techniques. Constraint-based methods build a knowledge base of libraries and the types contained within (Dong et al., 2022; Subramanian et al., 2014). The state-of-the-art technique SnR (Dong et al., 2022) leverages this knowledge base and uses constraints extracted from the code snippet to identify the precise types being used. In contrast, ML-based approaches (Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2024) learn from prior examples of type usage in existing code to perform type inference.
@@
ML-based techniques often produce incorrect inferences, due to their limited understanding of code structure and the rules governing Java’s type system. Recent advancements in large language models (LLMs), such as OpenAI’s proprietary GPT family and Meta’s open-weight Llama models, offer new possibilities for overcoming these limitations. By leveraging training on diverse and extensive datasets, LLMs have shown potential in performing various software engineering tasks (Hou et al., 2024). Prior studies (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024) suggest that LLMs achieved performance comparable to that of the state-of-the-art, SnR.
@@
However, the evaluation of the prior techniques (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024) is potentially affected by data leakage (Inan et al., 2021) where the training dataset includes the benchmark dataset, and data leakage allows LLMs to regurgitate the training data rather than conducting type inference. The benchmark suite, StatType-SO, used for evaluating type inference techniques has been publicly available on GitHub since 2017 (Sta, 2017), with the full suite available since 2023 (Sta, 2023). As shown in Table 1, recent state-of-the-art LLMs have knowledge cutoff dates in late 2023, creating the potential for leakage. Thus it remains uncertain whether the LLMs’ strong performance stems from their ability to understand the semantics of the code snippets or merely from retrieving the ground truth from their training data. Since the exact training data for LLMs are kept confidential, detecting such leakage is difficult (Carlini et al., 2021; Nasr et al., 2023). While recent work has called attention to this data leakage problem for software engineering research (Ozkaya, 2023; Sallou et al., 2024), no study has thoroughly, empirically evaluated LLMs’ performance on type inference on code snippets.
In this paper, we thoroughly evaluate LLMs’ type inference performance on Java code snippets to better understand the strengths and limitations of LLMs on type inference tasks, and uncover potential limitations of LLMs. We organized our inquiry into the following three research questions.
RQ1: How well do LLMs perform type inference on unseen code snippets?
@@
To address data leakage, we generated a new dataset named ThaliaType using Thalia, a program synthesis technique originally designed for testing static type systems. ThaliaType comprises 300 code snippets guaranteed to be unseen by the evaluated LLMs, thereby avoiding data leakage. Furthermore, we used the same set of libraries as used in StatType-SO to generate ThaliaType, to (1) enable a direct comparison of type inference performance between the two datasets, and (2) ensure the evaluated LLMs were consistently and sufficiently trained on the same set of types used in both datasets. By evaluating LLMs on ThaliaType and contrasting those results with those from StatType-SO, we can better understand the impact of data leakage on performance.
RQ2: To what extent do LLMs understand the semantics of code snippets?
@@
To investigate whether LLMs truly understand the semantics of code snippets during type inference, we designed semantic-preserving code transformations to create syntactically different versions of the input code (section 5). These transformations were intended to hinder LLMs’ reliance on memorized training data by making direct recall more challenging. By observing how LLMs’ performance changes on these semantically equivalent but syntactically distinct versions, we can gain insights into their ability to grasp the underlying semantic meaning of the code.
RQ3: What minimal syntax elements are sufficient for LLMs’ type inference?
Lastly, we sought to distill the minimal syntax elements sufficient for LLMs’ inference through delta debugging (Zeller and Hildebrandt, 2002) (section 6). By minimizing the input code to the LLMs while maintaining the LLMs’ inference results, we assess whether LLMs rely on superficial patterns from training data rather than genuinely analyzing code semantics. If LLMs are truly performing type inference rather than leveraging memorized patterns, essential elements such as simple names should be retained in the minimized snippets.
@@
Following the recommendation from prior software engineering research (Sallou et al., 2024), we evaluate LLM’s performance using two open-source models Llama3.1:8b and Llama3.1:70b, alongside state-of-the-art closed-source models GPT-4o and GPT-4o-mini (see Table 1(b)). As a baseline, we compare these models against SnR (Dong et al., 2022), the state-of-the-art constraint-based type inference technique. This evaluation aims to provide deeper insights into the potential limitations of using LLMs and to guide future research in evaluating LLMs’ performance on other software engineering tasks.
In RQ1, we found LLMs are highly likely to be affected by the data leakage issue and are prone to making incorrect inferences on unseen code snippets with up to 59% decrease in precision and up to 72% decrease in recall for Llama3.1:8b. On ThaliaType, even the best LLM GPT-4o only obtained a precision of 54.74% and a recall of 44.54% . On the other hand, SnR not affected by the data leakage issue, obtained a precision of 84.15% and a recall of 84.43% on ThaliaType. This stark discrepancy suggests LLMs are affected by data leakage, thus prompting further investigation into their type inference performance.
In RQ2, we found that LLMs performed significantly worse on transformed code snippets that were syntactically different but semantically equivalent. Applying all transformations caused a significant drop in F1 scores on the StatType-SO dataset across all LLMs. This suggests that LLMs rely heavily on the exact syntactic structure, indicating a lack of semantic understanding that limits their generalizability in type inference.
In RQ3, delta debugging revealed that LLMs can infer the same types using only 6.37% of the tokens in StatType-SO, with the number of unique simple names decreasing by 41.49% after minimization. Despite the type inference task primarily involves inferring FQNs for the types in the code snippet, LLMs would still correctly infer FQNs even when the types were absent. This suggests that StatType-SO may enable LLMs to perform well by exploiting superficial patterns in code snippets; this performance may be due to data leakage rather than a genuine understanding of code semantics.
Our experiments indicate that prior evaluations of LLMs were likely influenced by data leakage rather than a genuine comprehension of the semantics of the code snippets. To address these in the future, it is essential to evaluate LLMs on benchmarks that effectively measure their performance and are guaranteed to be excluded from the LLMs’ training data such as using Thalia to generate unseen code snippets.
@@
We proposed a new dataset ThaliaType along with an approach to creating such new datasets to rigorously evaluate type inference performance of LLMs while mitigating the potential biases introduced by data leakage. With the help of ThaliaType, we demonstrated that prior evaluations using StatType-SO are likely subjected to the data leakage problem and thus the performance results are inflated and misleading.
We investigated how well LLMs understand the semantics of code snippets by experimenting with whether semantic-preserving transformations on the code snippets affect the type inference performance of LLMs. The evaluation results on StatType-SO and ThaliaType show that the precision and recall of LLMs on type inference are dependent on the exact content in the code snippet despite preserving the semantic meaning of the code, indicating the LLMs lack a deep comprehension of the semantics and heavily rely on superficial syntactic patterns for type inference.
We further investigated the essence of LLMs performing type inference by isolating the syntactic elements sufficient for LLMs’ using delta debugging. The results show that some essential simple names are not required by LLMs for type inference especially on StatType-SO, indicating LLMs may rely more on exploiting superficial patterns from their training data than on the semantic information in code snippets.
This study provides a novel approach to rigorously assess LLMs’ comprehension of program semantics. To enable replication, reproduction, and further research on LLMs for code-related tasks, we have made our dataset and code publicly available at https://github.com/uw-pluverse/thalia-type.
@@ 2. Type Inference on Java Code Snippets
Type inference on code snippets aims to automatically determine the precise types within the code, specifically their FQNs (Subramanian et al., 2014; Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2024). This process is essential for effectively reusing code snippets from sources like StackOverflow, where type information is often incomplete or missing entirely.
2.1. Machine Learning-Based Type Inference
Machine learning-based methods use statistical models and learned representations of code to conduct type inference and resolve ambiguous simple names in code snippets (Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2024). Of these, LLMs have recently achieved state-of-the-art performance on type inference on the StatType-SO (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024) dataset. Existing approaches typically employ a simple but effective prompt to guide LLMs in generating FQNs for types in the code snippet (Huang et al., 2023a; Kabir et al., 2024; Chen et al., 2024). Notably, Kabir et al. (2024) extended the single-prompt approach by incorporating a secondary prompt to address compilation errors in the code snippet generated by the first prompt. This iterative process achieved a precision of 99.5% on the StatType-SO dataset. In contrast, our work focuses on the issue of data leakage in type inference rather than repairing code snippets. Data leakage is a fundamental problem that compromises the validity of results, regardless of whether a single-prompt or iterative approach is used. To ensure consistency and comparability with prior research, we adopted a single-prompt approach for type inference. Following established practices (Huang et al., 2023a; Kabir et al., 2024), we developed a straightforward but highly effective prompt for LLM-based type inference.
Prompt Template. Our prompt, shown in Figure 1, is designed to guide the LLM in inferring the type information in Java code snippets. It includes two key components: (1) a system message that sets the LLM’s role as a helpful assistant, and (2) a user instruction explicitly requesting the addition of import statements for Java code. The placeholder {input_code} is replaced with the actual code snippet during inference. Figure 1(a) presents the exact wording of the prompt, while Figure 1(b) demonstrates the corresponding output generated by the LLM for an example code snippet. Despite its simplicity, our prompt achieves high accuracy and aligns with best practices for LLM prompt engineering (LLM, 2024). Notably, our result using GPT-4o on the StatType-SO dataset represents a 7% improvement over the prior results (Kabir et al., 2024), demonstrating the robustness of our prompt while maintaining a similar level of complexity.
| System | You are a helpful programming assistant. |
| {input_code} |
2.2. Constraint-Based Type Inference with SnR
SnR (Dong et al., 2022) is the state-of-the-art technique that uses constraint-based type inference (Dong et al., 2022; Subramanian et al., 2014). It has proven highly effective for inferring types in code snippets from the StatType-SO dataset. To illustrate the information SnR leverages, consider the example code in Figure 2(a). SnR processes the snippet by extracting constraints, such as identifying a type with the simple name Logger, which has a getLogger method that accepts a java.lang.String argument. Using the knowledge base shown in Figure 2(c), SnR assigns java.util.logging.Logger to Logger, satisfying the constraints. This approach precisely eliminates incorrect candidates, such as org.slf4j.Logger, which lacks the required getLogger method.
@@ 2.2. Constraint-Based Type Inference with SnR
3. Datasets
We utilize two datasets StatType-SO and ThaliaType for evaluating LLMs’ type inference performance on code snippets.

### v2

# Unmasking the Genuine Type Inference Capabilities of LLMs for Java Code Snippets

@@ Abstract.
Type inference is a crucial task for reusing online code snippets. Although code snippets are prevalently shared on platforms like StackOverflow, they frequently lack essential type information, such as fully qualified names (FQNs) and required libraries. Recent studies have leveraged Large Language Models (LLMs) to perform type inference for such code snippets, demonstrating promising performance. However, these evaluations may suffer from data leakage, as the benchmark suite, StatType-SO, used for evaluation has been publicly available on GitHub since 2017. Consequently, it remains uncertain whether the strong performance of LLMs reflects genuine semantic understanding of code or is due to the ground truth being included in the training set.
This paper strives to comprehensively evaluate the genuine type inference capabilities of LLMs on Java code snippets and identify potential limitations of LLMs. First, we created ThaliaType, a new, previously unreleased benchmark suite designed for type inference evaluation. Second, using the StarCoder2 LLM as a baseline, we uncovered data leakage from StatType-SO in StarCoder2’s open-source training set and observed that other state-of-the-art LLMs exhibit similar performance drops when evaluated on ThaliaType, with precision decreasing by up to 59% and recall by up to 72%. Finally, we developed semantic-preserving code transformations to further investigate the capabilities of LLMs in understanding the execution semantics of code snippets. Our results showed that the performance of LLMs on StatType-SO is far less robust to these transformations than on ThaliaType, suggesting that the performance on StatType-SO may be biased by data leakage and have limited generalizability.
These findings highlight the importance of carefully designed, leakage-free benchmarks for evaluating LLMs on type inference tasks. We recommend future studies adopt ThaliaType to ensure rigorous and reliable assessments of LLMs’ genuine type inference capabilities.
@@ 1. Introduction
Type inference for code snippets can assist developers in recovering missing type information needed to reuse these code snippets (Dong et al., 2022; Subramanian et al., 2014; Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023b; Kabir et al., 2025). For example, current approaches to type inference for Java code snippets generally fall into two categories: constraint-based and machine learning (ML)-based techniques. Constraint-based methods build a knowledge base of libraries and the types contained within (Dong et al., 2022; Subramanian et al., 2014). The state-of-the-art technique SnR (Dong et al., 2022) leverages this knowledge base and uses constraints extracted from the code snippet to identify the precise types being used.
@@
ML-based approaches (Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2025) learn from prior examples of type usage in existing code to perform type inference. They often produce incorrect inferences due to their limited understanding of code structure and the rules governing Java’s type system. Recent advancements in large language models (LLMs), such as OpenAI’s proprietary GPT family and Meta’s open-weight Llama models, offer new possibilities for overcoming these limitations. By leveraging training on diverse and extensive datasets, LLMs have shown potential in performing various software engineering tasks (Hou et al., 2024). Prior studies (Huang et al., 2023a; Kabir et al., 2025; Chen et al., 2024) suggest that LLMs achieved performance comparable to that of the state-of-the-art SnR.
| Benchmark Suite | First Available | #Code Snippets |
| StatType-SO | 2017 (see Table 2) | 267 |
| ThaliaType | Always New | 300 |
| Model | Version | Knowledge Cutoff | Data Leakage |
| StarCoder2:15b | instruct-v0.1-q4_0 | September 2023 (Lozhkov et al., 2024) | ✓ |
| GPT-4o | 2024-08-06 | October 2023 (OPE, 2024a) | ? |
| GPT-4o-mini | 2024-07-18 | October 2023 (OPE, 2024a) | ? |
| Llama3.1:8b | instruct_q4_0 | December 2023 (Meta, 2024) | ? |
| Llama3.1:70b | instruct_q4_0 | December 2023 (Meta, 2024) | ? |
@@
However, the evaluation of these prior studies (Huang et al., 2023a; Kabir et al., 2025; Chen et al., 2024) is potentially affected by data leakage (Inan et al., 2021) where the training dataset includes the benchmark dataset together with the ground truth, and data leakage allows LLMs to regurgitate the training data rather than conducting type inference. The benchmark suite StatType-SO, used for evaluating type inference techniques has been fully, publicly available on GitHub since 2017 (Sta, 2017). As shown in Table 1, recent state-of-the-art LLMs have knowledge cutoff dates in late 2023, creating the potential for leakage. Thus, it remains uncertain whether the LLMs’ strong performance stems from their ability to understand the semantics of the code snippets or merely from retrieving the ground truth from their training data. While recent work has called attention to this data leakage problem for software engineering research (Ozkaya, 2023; Sallou et al., 2024), no study has thoroughly, empirically evaluated LLMs’ performance on type inference on code snippets. Because prior work relied on GPT models where the exact training data is kept confidential, detecting such leakage is difficult (Carlini et al., 2021; Nasr et al., 2023).
To address this limitation, we first investigated The Stack v2 (BigCode, 2024) (simply referred to as StackV2), an open-source dataset used to train the StarCoder2 model. Our exploration revealed that all code snippets from StatType-SO with ground truth were present in StackV2. This means that StarCoder2 was trained using code snippets along with their expected import statements from StatType-SO. Thus, StarCoder2’s performance is affected by data leakage, and can potentially recall the code snippets and their expected import statements from training rather than performing type inference by analyzing the semantics of the code snippets. To better understand the behavior of LLMs, we compare StarCoder2’s performance against Llama and GPT models. We organized our inquiry into the following two research questions as illustrated in Figure 1.
In this paper, we thoroughly evaluate LLMs’ type inference performance on Java code snippets, focusing on their ability to leverage type information, the impact of data leakage, and their generalizability to unseen code.
@@
RQ1: How well do LLMs perform type inference on unseen code snippets? To address data leakage, and investigate LLMs’ genuine ability to perform type inference using the semantic information present in code snippets, we generated a new benchmark suite named ThaliaType using Thalia, a program synthesis technique originally designed for testing static type systems (Sotiropoulos et al., 2024). ThaliaType comprises 300 code snippets that are guaranteed to be unseen by the evaluated LLMs, thereby avoiding data leakage. Unlike StatType-SO, which was published in 2017 (Table 1(a)), new code snippets can always be generated using our replication package, enabling future evaluations without data leakage. To generate ThaliaType, we used the same set of libraries used in StatType-SO to (1) enable a direct comparison of type inference performance between the two benchmark suites and (2) ensure the evaluated LLMs were consistently and sufficiently exposed to the types used in both benchmark suites during training. By evaluating LLMs on ThaliaType and contrasting the results with those from StatType-SO, we can gain a better understanding of the impact of data leakage on performance.
@@
RQ2: To what extent do LLMs understand the execution semantics of code snippets? To investigate whether LLMs truly understand the execution semantics of code snippets during type inference, we designed semantic-preserving code transformations to create syntactically different versions of the input code (section 6). These transformations were intended to hinder LLMs’ reliance on memorized training data by making direct recall more challenging. By observing how LLM’ performance changes on these semantically equivalent but syntactically distinct versions, we can gain insights into their ability to grasp the underlying semantic meaning of the code.
@@
Following recommendations from prior software engineering research (Sallou et al., 2024), we evaluate LLM’s performance using three open-weight models StarCoder2, Llama3.1:8b, and Llama3.1:70b, along with state-of-the-art closed-source models GPT-4o and GPT-4o-mini (see Table 1(b)). Among these, StarCoder2 serves as a baseline model with confirmed data leakage (see section 4). For additional comparison, we include SnR (Dong et al., 2022), a state-of-the-art constraint-based type inference technique, as a non-LLM baseline. This evaluation aims to provide deeper insights into the potential limitations of using LLMs and to guide future research evaluating LLMs’ performance on other software engineering tasks.
Findings of RQ1. In RQ1, we found that LLMs are prone to making incorrect inferences on code snippets from ThaliaType, likely because these snippets are truly unseen and not affected by data leakage. All tested LLMs exhibited a similar decline in performance as StarCoder2:15b which experienced a 66.8% decrease in F1-score. For example, Llama3.1:8b and GPT-4o showed comparable declines of 67.2% and 48.5% respectively. On both StatType-SO and ThaliaType, StarCoder2:15b outperformed Llama3.1:8b but was outperformed by Llama3.1:70b, GPT-4o-mini, and GPT-4o. In contrast, SnR, which is not affected by data leakage, obtained a precision of 84.15% and a recall of 84.43% on ThaliaType. The consistent drop in LLMs’ performance between the baseline StarCoder2:15b and other models suggests that all tested LLMs are potentially affected by data leakage. In addition, although StarCoder2:15b was trained on StatType-SO snippets, it only achieved an F1-score of 81.67% , indicating that even in the presence of data leakage, models do not perfectly recall the training data.
Findings of RQ2. In RQ2, LLMs showed robustness to individual transformations on both benchmark suites, and also to the combined transformation on ThaliaType code snippets. StarCoder2:15b showed only a 3.0% decrease in F1-score in the worst case, while Llama3.1:8b experienced a 7.9% drop. However, combined transformations on StatType-SO consistently led to performance degradation across all models, including a 16.8% reduction for StarCoder2 and up to 30.0% for Llama3.1:8b. Notably, these same combined transformations did not consistently degrade performance on ThaliaType. In fact, StarCoder2:15b experienced a slight improvement in F1-score. One possible explanation is the presence of data leakage, which may allow models to recall answers from training. When transformations introduce enough variation to prevent direct recall, the models are forced to rely on semantic reasoning. Under these conditions, their performance on StatType-SO aligned more closely with that on truly unseen data.
Our results suggest that future evaluations of LLM-based techniques should also assess generalizability, defined as the ability for LLM-based techniques to maintain consistent performance across unseen code snippets. This can be achieved by incorporating both transformed code snippets and ThaliaType, rather than relying solely on StatType-SO, which may suffer from data leakage.
@@
We introduce ThaliaType, a new benchmark suite constructed through a novel adaptation of Thalia for rigorously evaluating the type inference performance of LLMs while mitigating potential bias from data leakage. Using ThaliaType, we demonstrate that prior evaluations based on StatType-SO are likely affected by data leakage, leading to inflated and potentially misleading performance results.
We evaluate whether LLMs understand the execution semantics of code snippets by applying transformations that preserve the semantics and comparing their effects. Our results on StatType-SO and ThaliaType show that LLMs are generally able to maintain performance across individual transformations. However, complex transformations on StatType-SO result in disproportionately large performance drops compared to both their effects on ThaliaType and the individual transformations. This lack of generalizability can be caused by data leakage where models learned the correct answer from training.
To enable generation of new code snippets for future research on LLM type inference for code snippets, we have made our benchmark suite and code publicly available at https://github.com/uw-pluverse/thalia-type.
@@ 2. Type Inference on Java Code Snippets
Type inference on code snippets aims to automatically determine the precise types within the code, specifically their FQNs (Subramanian et al., 2014; Dong et al., 2022; Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2025). This process is essential for effectively reusing code snippets from sources like StackOverflow, where type information is often incomplete or missing entirely.
@@
2.1. Machine Learning-Based Type Inference
Machine learning-based methods use statistical models and learned representations of code to conduct type inference and resolve ambiguous simple names in code snippets (Phan et al., 2018; Saifullah et al., 2020; Huang et al., 2023a; Kabir et al., 2025). Of these, LLMs have recently achieved state-of-the-art performance on type inference on the StatType-SO (Huang et al., 2023a; Kabir et al., 2025; Chen et al., 2024) benchmark suite. Existing approaches typically employ a simple but effective prompt to guide LLMs in generating FQNs for types in the code snippet (Huang et al., 2023a; Kabir et al., 2025; Chen et al., 2024). Notably, Kabir et al. (2025) extended the single-prompt approach by incorporating a secondary prompt to address compilation errors in the code snippet generated by the first prompt. This iterative process achieved 99.5% precision on StatType-SO. In contrast, our work focuses on the issue of data leakage in type inference rather than repairing code snippets. Data leakage is a fundamental problem that compromises the validity of results, regardless of whether a single-prompt or iterative approach is used. To ensure consistency and comparability with prior research, we adopted a single-prompt approach for type inference. Following established practices (Huang et al., 2023a; Kabir et al., 2025), we developed a straightforward but highly effective prompt for LLM-based type inference.
Prompt Template. Our prompt, shown in Figure 3, is designed to guide the LLM in inferring the type information in Java code snippets. It includes two key components: (1) a system message that sets the LLM’s role as a helpful assistant, and (2) a user instruction explicitly requesting the addition of import statements for Java code. The placeholder {input_code} is replaced with the actual code snippet during inference. Figure 3(a) presents the exact wording of the prompt, while Figure 3(b) demonstrates the corresponding output generated by the LLM for an example code snippet. Despite its simplicity, our prompt achieves high accuracy and aligns with best practices for LLM prompt engineering (LLM, 2024). Notably, our results using GPT-4o on the StatType-SO benchmark suite represent a 7% improvement over the prior results (Kabir et al., 2025), demonstrating the robustness of our prompt while maintaining a similar level of complexity.
| System | You are a helpful programming assistant. |
| {input_code} |
2.2. Constraint-Based Type Inference with SnR
SnR (Dong et al., 2022) is the state-of-the-art technique that uses constraint-based type inference (Dong et al., 2022; Subramanian et al., 2014). It has proven highly effective for inferring types in code snippets from StatType-SO. To illustrate the information SnR leverages, consider the example code in Figure 2(a). SnR processes the snippet by extracting constraints, such as identifying a type with the simple name Logger, which has a getLogger method that accepts a java.lang.String argument. Using the knowledge base shown in Figure 2(c), SnR assigns java.util.logging.Logger to Logger, satisfying the constraints. This approach precisely eliminates incorrect candidates, such as org.slf4j.Logger, which lacks the required getLogger method.
3. Benchmark Suites for Evaluating Type Inference
We utilize two benchmark suites StatType-SO and ThaliaType for evaluating LLMs’ type inference performance on code snippets.